OpenStudy (anonymous):

EVALUATE THE EXPONENTIAL EXPRESSION BELOW 3*(5+4 )^2 -4^2 = ?? HELP PLEASE THANK U

jhonyy9 (jhonyy9):

3*81-16 =243-16=227

jhonyy9 (jhonyy9):

ok ?